OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDALS!!!!! Solve, using the substitution method. d + e = 2 d – e = 4 The solution is (3, –1). The solution is (5, 1). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

There is no solution??

zepdrix (zepdrix):

Let's look at the second equation:\[\Large\rm d-e=4\]If we add \(\Large\rm e\) to each side we get,\[\Large\rm d=4+e\]I'm going to color it like this,\[\Large\rm \color{orangered}{d=4+e}\]And maybe it will help clarify what we're about to do.

zepdrix (zepdrix):

Sorry had to fix my post >.<

zepdrix (zepdrix):

We're going to use this and fix in some stuff in our first equation:\[\Large\rm \color{orangered}{d} + e = 2\]

zepdrix (zepdrix):

From working with our first equation, we found that the letter d can be replaced by \(\Large\rm 4+e\).

zepdrix (zepdrix):

That's the whole idea with substitution. We want to replace one of the letters with some stuff involving the other letter.

zepdrix (zepdrix):

So our equation:\[\Large\rm \color{orangered}{d} + e = 2\]becomes,\[\Large\rm \color{orangered}{4+e} + e = 2\]

zepdrix (zepdrix):

Do you understand how to solve for the value of \(\Large\rm e\) from this point?

OpenStudy (anonymous):

you know I appreciate you trying to help me learn this but honestly I really need this answer so could you please tell me which answer is correct?

zepdrix (zepdrix):

No.

OpenStudy (agl202):

d + e = 2 de = 4 Solve for e for the first equation. e = -d + 2 Substitute that for the second equation and solve for d: d(-d + 2) = 4 -d² + 2d = 4 -d² + 2d - 4 = 0 By quadratic formula, which is d = (-b² ± √(b² - 4ac))/2a: a = -1 b = 2 c = -4 Substituting those values to the formula gives: d = (-2 ± √(2² - 4(-1)(-4)))/2(-1) d = (-2 ± √(4 - 16))/-2 d = (-2 ± √-12)/-2 Since the discriminant b² - 4ac gives nonreal solutions, there are NO REAL SOLUTIONS for the system of equations.

OpenStudy (agl202):

So correct answer is no real solution!!! Sorry have no clue why there are question marks up there!:(

OpenStudy (anonymous):

Thank you so much could you help me with more? it'd really help

OpenStudy (agl202):

sure...

zepdrix (zepdrix):

The question didn't involve de=4. I can't figure out where you got that from...

OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 4x – 6 8x – 2y = 14 (14, 12) (12, 14) There are an infinite number of solutions. There is no solution.

OpenStudy (anonymous):

@Agl202

OpenStudy (agl202):

Ok so, The lines are parallel so they do not intersect. Meaning no solutions.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Solve, using the substitution method. 2x + 3y = 1 y = x – 8 (0, –8) (2, –6) (4, –4) (5, –3)

OpenStudy (agl202):

@zepdrix You just gonna do the math correctly, then u will understand were I got it from! Ok?

OpenStudy (agl202):

Can u put ur last question into a new one? Would realy help me... Thx: )

zepdrix (zepdrix):

Loooool, but you didn't do the math correctly >.< You chose the wrong answer...

OpenStudy (agl202):

@zepdrix Sorry I have proof!!!! Ok?